OpenStudy (anonymous):

find the derivative: (square rt(2))/(3y^3)

OpenStudy (anonymous):

is this \[\frac{\sqrt{2}}{3y^3}\]?

OpenStudy (anonymous):

yes sir thats it! now it says to find the derivative

OpenStudy (anonymous):

ok so suppose the square root of two was not there and you wanted \[\frac{d}{dy}\frac{1}{3y^3}\] could you find that?

OpenStudy (anonymous):

i know you make it (3y^3)^-1

OpenStudy (anonymous):

or better still how about \[\frac{d}{dy}\frac{1}{y^3}\]

OpenStudy (anonymous):

i would not suggest doing it that way. first let's find \[\frac{d}{dy}\frac{1}{y^3}\]

OpenStudy (anonymous):

for this one you might want to write \[\frac{1}{y^3}=y^{-3}\] so \[\frac{d}{dy}\frac{1}{y^3}=\frac{d}{dy}y^{-3}=-3y^{-4}=\frac{-3}{y^4}\]

OpenStudy (anonymous):

then the coefficient \[\frac{\sqrt{2}}{3}\] gets multiplied by the result. so you get \[\frac{\sqrt{2}}{3}\times \frac{-3}{y^4}\] which you can clean up some

OpenStudy (anonymous):

ok ok ok awesome thank you very much

OpenStudy (anonymous):

yw. don't let the constants fool you. take them out side of the derivative and put them in at the end

OpenStudy (anonymous):

ok my prof at Texas A&M says to never deal with rational exponents so she says to make them negative and change them back at the end but your way is easier